OpenStudy (anonymous):

Two forces F1 and F2 act on a 5.00 kg object. Taking F1=20.0N and F2=15.00N, find the acceleration of the object for the configurations: a) mass m with F1 acting in the positive x direction and F2 acting perpendicular in the positive y direction b) mass m with F1 acting in the positive x direction and F2 acting on the object at 60 degrees above the horizontal.

OpenStudy (anonymous):

Fr=sqrt(20^2+15^2)= 25N a=Fr/m=25N/5kg=5m/s^2

OpenStudy (anonymous):

first find the resultant force which has equivalent effect as the individual forces applied separately

OpenStudy (anonymous):

since the two forces are perpendicular to each other you can use pythagorus theorem. the resulting force is the force that i designated as Fr

OpenStudy (anonymous):

once you get the force divide by the mass of the object.